Write a Pythagorean triplet whose one member is. 
18

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

For any natural number m, we know that 2m, m2–1, m2+1 is a Pythagorean triplet.
2m=18
⇒m=18/2
=9
m2–1=92–1
= 81–1
= 80
m2+ 1 = 92+ 1
= 81+1
=82
∴ (18, 80, 82) is a Pythagorean triplet

Thanks